# CPRet: A Dataset, Benchmark, and Model for Retrieval in Competitive Programming

### v1

@@ 3.1 Overview of Retrieval Tasks
Table 1: Statistics of the four retrieval tasks in CPRet, including the number of training and test items. $L_{\text{Query}}$ and $L_{\text{Corpus}}$ denote average token lengths.
@@ 3.3.1 CPRetriever-Code: Multi-Positive Contrastive Learning for Problem-Code Alignment
This group-based formulation improves representation quality by explicitly modeling the structure of multiple correct outputs, leading to more stable and discriminative embeddings for code-centric tasks.
@@ 4.1 Experimental Setup
We develop two retrieval models based on the SFR-Embedding-Code-2B_Rsfrcode embedding model, which is built on top of Gemma-2-2Bgemma and further pre-trained on code-related tasks. Detailed training configurations are provided in Appendix A.1.
@@ 4.2 Main Results
Medium multilingual-e5-large me5 general 560M 4.27 18.51 19.19 65.02 26.75
@@
gte-Qwen2-1.5B-inst. gte general 1.5B 10.76 23.41 27.06 69.15 32.60
@@
CPRetreiver-Code code 2B 70.40 70.59 38.68 81.45 65.28
CPRetreiver-Prob code 2B 56.50 70.68 60.06 90.74 69.50
@@ A.1 Additional Experimental Setup Details
For our contrastive pretraining, we use a batch size of 1024, enabled by gradient caching[41] to optimize GPU memory usage. The model is trained with a learning rate of $3\times 10^{-8}$, temperature $\tau=0.07$, and weight decay of 0.01, using the AdamW[42] optimizer. Each query is paired with $m=16$ positive samples, randomly sampled from the dataset. If fewer than $m$ positive samples are available, duplicates are used to pad the group. The model is trained for 20 epochs on 8 NVIDIA A800 GPUs, taking approximately 70 hours to complete. This stage focuses primarily on problem-code representation alignment. The resulting model, which demonstrates strong performance on code-related retrieval tasks, is denoted as CP-Retriever-Code.
@@
A.2 Ablation study
A.2.1 Ablation study on Group-infoNCE
@@
Table 4 presents the performance of different loss functions and training settings on the Text-to-Code and Code-to-Code retrieval tasks. Here, InfoNCE corresponds to Equation 1, InfoNCE (multi-pos) refers to the multi-positive version in Equation 2, and Group-InfoNCE is our proposed method defined in Equation 3.
We make the following key observations:
  • Data augmentation consistently improves performance across comparable configurations, demonstrating its effectiveness.
  • InfoNCE (multi-pos) underperforms compared to the standard single-positive InfoNCE, suggesting that naïvely adding multiple positives may not be beneficial.
  • Increasing the number of positives $m$ generally leads to better performance, indicating that models benefit from more positive examples.
The ablation results confirm that our proposed Group-InfoNCE loss leads to substantial improvements in both Text-to-Code and Code-to-Code tasks, validating its effectiveness for joint representation learning of problems and solutions.
A.2.2 Ablation Study on Fine-Tuning Data Composition
@@
A.3 Further Details on the CPRet-PCPCD
A.3.1 Language and Problem Format Statistics
@@
A.3.2 Programming Language Distribution
@@
A.4 Data Source Details
@@
A.4.1 Competitive Programming Problems and Codes Data(CPRet-PCPCD)
@@
A.4.2 Duplicate Programming Problem Pairs
@@
A.4.3 Simplified and Full Problem Description Pairs
@@
A.5 Extended Results: Model Performance across Similarity Bins
@@
A.6 Cases of Retrieval Tasks
A.6.1 Text-to-Code and Code-to-Code
@@
https://atcoder.jp/contests/arc162/tasks/arc162_a Problem Statement There are $N$ people, numbered from $1$ to $N$, who participated in a round-trip race between two points. The following information is recorded about this race. The outward times of any two people were different, and person $i$ $(1\leq i\leq N)$ had the $i$-th fastest outward time. The round-trip times (the sum of the outward and return times) of any two people were different, and person $i$ $(1\leq i\leq N)$ had the $P_{i}$-th fastest round-trip time. The person (or persons) with the fastest return time was awarded the fastest return award. Here, $P_{1},P_{2},\dots,P_{N}$ is a permutation of $1,2,\dots,N$. How many people could have received the fastest return award? There are $T$ test cases. Answer each of them. Constraints $1\leq T\leq 500$ $2\leq N\leq 10^{3}$ $P_{1},P_{2},\dots,P_{N}$ is a permutation of $1,2,\dots,N$ All input values are integers The sum of $N$ over all test cases does not exceed $10^{3}$ Input Format The input is given from standard input in the following format:
@@ A.6.2 Problem-to-Duplicate
https://codeforces.com/contest/1702/problem/G2 Problem: G2. Passable Paths (hard version) You are given a tree with $n$ vertices numbered from $1$ to $n$. A tree is an undirected connected graph with $n-1$ edges and no cycles. You are also given $q$ queries. In each query, you are given a set of $k$ vertices. For each query, determine whether there exists a simple path in the tree that passes through all the given $k$ vertices. Input The first line contains an integer $n$ ($2\leq n\leq 2\times 10^{5}$) — the number of vertices in the tree. Each of the next $n-1$ lines contains two integers $u$ and $v$ ($1\leq u,v\leq n$) — the endpoints of an edge in the tree. The next line contains an integer $q$ ($1\leq q\leq 2\times 10^{5}$) — the number of queries. Each of the next $q$ lines describes a query: The first integer $k$ ($2\leq k\leq n$) — the number of vertices in the query. Followed by $k$ integers $v_{1},v_{2},\ldots,v_{k}$ ($1\leq v_{i}\leq n$) — the vertices in the query. Output For each query, print YES if there exists a simple path that passes through all the given $k$ vertices. Otherwise, print NO.
https://www.codechef.com/JULY21A/problems/KPATHQRY Problem: KPATHQRY — Path Queries on Trees You’re given a tree with $N$ vertices numbered from $1$ to $N$. Your goal is to handle $Q$ queries. For each query, you are given $K$ nodes $v_{1},v_{2},\ldots,v_{K}$. Find whether there exists a simple path in the tree that covers all the given vertices. Input The first line contains a single integer $T$ — the number of test cases. For each test case: The first line contains a single integer $N$ — the number of vertices. Each of the following $N-1$ lines contains two integers $u$ and $v$ — an edge in the tree. The next line contains a single integer $Q$ — the number of queries. Each of the following $Q$ lines describes a query: * Starts with $K_{i}$ — the number of vertices in the query. * Followed by $K_{i}$ integers: $v_{1},v_{2},\ldots,v_{K_{i}}$. Output For each query, print "YES" if a simple path covering all the given nodes exists, otherwise print "NO". You may print each character of the string in uppercase or lowercase (for example, "yEs", "yes", "Yes", and "YES" will all be treated as identical). Constraints $1\leq T\leq 10$ $1\leq N\leq 10^{5}$ $1\leq u,v,v_{j}\leq N$ $1\leq Q\leq 10^{5}$ $1\leq K_{i}\leq N$ The edges form a valid tree. All vertices in a single query are distinct. The sum of $N$ over all test cases does not exceed $2\cdot 10^{5}$. The sum of $K_{i}$ over all queries in a single test case does not exceed $10^{5}$. Subtasks Subtask #1 (100 points): Original constraints.
A.6.3 Simplified-to-Full

### v2

@@ 1 Introduction
Competitive programming problem retrieval poses unique challenges compared to standard code retrieval: problem statements are often abstract, narrative-driven, and may admit multiple fundamentally different solutions, requiring models to capture high-level algorithmic ideas rather than surface code patterns. This motivates the need for specialized problem-level retrieval models.
@@ 2.2 Embedding Models for Retrieval
High-quality retrieval has been shown to benefit downstream tasks like code generation via Retrieval-Augmented Generation (RAG), where an LLM leverages retrieved problem-solution pairs as context. For example, Shi et al. shi2024can explored solving Olympiad-level problems with RAG, Li et al. li2024can proposed critic-guided retrieval-augmented planning, and Mapcoder islam2024mapcoder introduced MapCoder, a multi-agent code generation framework using retrieval. These studies suggest that problem-level retrieval can substantially enhance LLM performance, motivating specialized embeddings for competitive programming.
While temporal splits and duplicate-problem detection are not new concepts—having been explored in other domains such as finance de2018advances and software community question answering hazra2023duplicate —their systematic application to competitive programming remains underexplored. Existing benchmarks (e.g., MTEBMTEB , CoIRCoIR ) often lack strict temporal separation, potentially inflating evaluation results. Our work extends these established principles to programming contests, where duplicate or temporally overlapping problems introduce unique challenges: algorithmically similar tasks may differ superficially in text but still bias performance. By constructing a benchmark with rigorous temporal splits and fine-grained similarity filtering, we aim to build fairer evaluation standards and reveal new insights into the relationship between problem similarity and model discrimination ability.
@@ 3.1 Overview of Retrieval Tasks
Table 1: Statistics of the four retrieval tasks in CPRet. We report the number of training and test items, where #Train-Code indicates the number of distinct solution codes, and #Train-Pair refers to the number of (anchor, positive) pairs used for training. $L_{\text{Query}}$ and $L_{\text{Corpus}}$ denote average token lengths.
@@ 3.3.1 CPRetriever-Code: Multi-Positive Contrastive Learning for Problem-Code Alignment
This group-based formulation improves representation quality by explicitly modeling the structure of multiple correct solutions, resulting in more stable and discriminative embeddings for code-centric tasks. Although Group-InfoNCE enforces consistency between a problem embedding and its set of correct solutions, it does not constrain the diversity of the solutions themselves. The variance-based regularization ensures that each solution maintains a consistent similarity to the problem embedding, effectively "surrounding" it in the code space without collapsing distinct approaches. In practice, these similarities converge around an average value, allowing the problem representation to capture the essential idea while accommodating diverse algorithmic strategies. Empirical results on both base models (see Tables 4 and 5) show consistent improvements in retrieval performance and robust, discriminative embeddings.
@@ 4.1 Experimental Setup
We develop two retrieval models based on the SFR-Embedding-Code-2B_Rsfrcode embedding model, which is built on top of Gemma-2-2Bgemma and further pre-trained on code-related tasks. We further evaluate our method on the Qwen3-Embedding-4Bqwen3emb model, released in June 2025, which demonstrates stronger retrieval capabilities. Detailed training configurations are provided in Appendix A.1.
@@ 4.2 Main Results
multilingual-e5-large me5 general 560M 4.27 18.51 19.19 65.02 26.75
@@
Qwen3-Embedding-0.6B qwen3emb general 600M 48.96 60.49 36.26 81.63 56.83
Medium gte-Qwen2-1.5B-inst. gte general 1.5B 10.76 23.41 27.06 69.15 32.60
@@
CPRetriever-Code code 2B 70.40 70.59 38.68 81.45 65.28
CPRetriever-Prob code 2B 56.50 70.68 60.06 90.74 69.50
Qwen3-Embedding-4B qwen3emb general 4B 66.62 71.97 56.59 89.39 71.15
CPRetriever-Code-Qwen3-4B code 4B 86.22 86.70 41.14 88.10 75.54
CPRetriever-Prob-Qwen3-4B code 4B 80.84 87.10 74.33 96.15 84.60
@@ 5 Limitations and Future Work
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62402264), and by the JC STEM Lab of AI for Science and Engineering, funded by The Hong Kong Jockey Club Charities Trust and the Research Grants Council of Hong Kong (Project No. CUHK14213224).
@@ A.1 Additional Experimental Setup Details
For our contrastive pretraining, we use a batch size of 1024, enabled by gradient caching[47] to optimize GPU memory usage. The model is trained with a learning rate of $3\times 10^{-8}$, temperature $\tau=0.07$, and weight decay of 0.01, using the AdamW[48] optimizer. Each query is paired with $m=16$ positive samples, randomly sampled from the dataset. If fewer than $m$ positive samples are available, duplicates are used to pad the group. The maximum input length is set to 1024 tokens. The model is trained for 20 epochs on 8 NVIDIA A800 GPUs, taking approximately 70 hours to complete. This stage focuses primarily on problem-code representation alignment. The resulting model, which demonstrates strong performance on code-related retrieval tasks, is denoted as CP-Retriever-Code.
@@
For experiments conducted with the Qwen3-Embedding-4B model, we adjusted several hyperparameters to better accommodate its larger architecture. Specifically, the maximum input length was increased to 2048 tokens, the contrastive learning temperature was set to $\tau=0.05$ (compared to the model’s default of 0.01), and the learning rate was set to $1\times 10^{-6}$. The total training time scaled approximately proportionally with the increase in model parameters.
A.1.1 Ablation Study on Group-InfoNCE
@@
Table 4 reports the performance of different loss functions and configurations on the Text-to-Code and Code-to-Code retrieval tasks. Here, InfoNCE corresponds to Equation 1, InfoNCE (multi-pos) refers to its multi-positive extension (Equation 2), and Group-InfoNCE denotes our proposed loss (Equation 3).
We observe the following:
  • Data augmentation consistently improves performance across comparable configurations, highlighting its robustness.
  • Naïvely increasing positive samples in InfoNCE (multi-pos) does not guarantee gains, suggesting that grouping structure is essential for effective multi-positive learning.
  • Larger $m$ values in Group-InfoNCE generally lead to higher accuracy, indicating that the model benefits from richer positive context.
Table 5: Multi-task training validation using Qwen3-4B-Embedding.
m Problem-Level Finetuning T2C C2C P2Dup S2Full Avg
No No 62.68 65.15 57.23 89.31 68.59
1 No 67.94 66.31 56.21 90.23 70.17
4 No 68.30 66.20 56.29 90.24 70.26
16 No 68.54 66.49 56.43 90.07 70.38
16 Yes 65.85 70.18 71.58 95.02 75.66
We further validated the effectiveness of Group-InfoNCE by repeating multi-task training with the Qwen3-4B-Embedding model using the same hyperparameter settings as in our main experiments (Table 5). Results show that fine-tuning on problem-level tasks slightly decreases Text-to-Code performance (from 68.54 to 65.85) but substantially improves problem-level tasks such as Problem-to-Duplicate and Simplified-to-Full. This consistent, minor trade-off supports our hypothesis that the embedding requirements of code-level and problem-level tasks are inherently different. Given this intrinsic conflict, we consider this small drop reasonable and do not pursue extensive hyperparameter optimization that might harm generalizability.
Overall, both ablation and multi-task analyses demonstrate that Group-InfoNCE effectively balances representation learning across heterogeneous tasks while maintaining strong cross-domain retrieval performance.
A.1.2 Ablation Study on Fine-Tuning Data Composition
@@
A.2 Supplementary Analysis for Figure 2
Figure 2 in the main paper presents two complementary perspectives on model performance trends. The left sub-plot focuses on illustrating the overall improvement of embedding models over time, highlighting representative models that achieved state-of-the-art (SOTA) performance at their respective publication dates. In contrast, the right sub-plot was designed to examine potential data leakage effects by including more recent, high-performing models that are more likely to exhibit such phenomena. As a result, the model sets in the two sub-figures are not identical.
To provide a more comprehensive view, Table 7 below reports the year-wise retrieval performance of additional models (e.g., CodeSage-large, Contriever-MSMarco, and E5-Base-v2) that were not included in the right sub-plot of Figure 2. Despite having lower absolute performance compared to the models shown in the right plot, these models display a consistent “decline-then-recovery” trajectory: strong performance in earlier years (e.g., 2015–2016), followed by a sharp degradation (2017–2019), and a partial recovery or stabilization in later years (2020–2024).
This trend mirrors the temporal performance decay pattern attributed to data leakage, as discussed in Section 3.2.1. Specifically, the observed recovery phase suggests that recently released problems are less contaminated by training data, leading to more realistic evaluations. Therefore, these supplementary results reinforce our main claim that data leakage can substantially inflate retrieval metrics in temporally overlapping benchmarks, and that rigorous temporal separation is critical for fair assessment.
Table 7: Supplementary results supporting Figure 2. Retrieval performance (%) of representative models across problem release years. The “decline-then-recovery” pattern indicates potential data leakage in earlier problems.
Model 2015 2016 2017 2018 2019 2020 2021 2022 2023 2024
SFR-Emb.-Code-2B 55.38 57.67 48.79 44.78 42.55 41.59 37.42 34.38 34.08 34.78
GTE-ModernBERT-Base 42.10 45.92 28.44 15.45 10.34 12.68 9.44 9.82 9.31 9.98
CodeSage-Large 40.65 43.57 27.71 14.67 9.10 11.76 9.33 7.67 7.21 8.08
Contriever-MSMarco 16.03 19.85 10.18 3.72 1.62 1.98 0.95 0.88 0.79 0.91
E5-Base-v2 22.35 26.15 14.02 5.16 2.27 2.59 1.69 1.32 1.18 1.29
A.3 Evaluation of Retrieval for RAG-based Competitive Programming
To investigate the impact of problem retrieval on downstream code generation tasks, we conducted experiments using Retrieval-Augmented Generation (RAG) in competitive programming problem-solving. We used problems from LiveCodeBench as the test set. To ensure temporal fairness, the retrieval pool was restricted to problems in our dataset published before April 2023, while LiveCodeBench problems appeared starting May 2023.
For each test problem, we retrieved the top-K most similar problems from the historical subset and provided their descriptions along with one accepted solution as context to the LLM. Three retrieval settings were compared: no retrieval (baseline), using Qwen3-Embedding-4B as the retriever, and using our CPRetriever-Prob (based on Qwen3-Embedding-4B).
For the code generation evaluation, we used three models from the Qwen-Coder series:
  • qwen-coder-turbo and qwen-coder-plus: non-reasoning models,
  • qwen3-coder-plus: a reasoning-capable model.
The results are summarized in Table 8.
Table 8: Pass@1 performance for RAG-based code generation with different retrievers.
Model Type RAG Model Pass@1
qwen-coder-turbo Non-reasoning None 0.3575
qwen-coder-turbo Non-reasoning Qwen3-Embedding-4B 0.3825
qwen-coder-turbo Non-reasoning CPRetriever-Prob 0.4075
qwen-coder-plus Non-reasoning None 0.5125
qwen-coder-plus Non-reasoning Qwen3-Embedding-4B 0.5275
qwen-coder-plus Non-reasoning CPRetriever-Prob 0.5500
qwen3-coder-plus Reasoning None 0.7675
qwen3-coder-plus Reasoning Qwen3-Embedding-4B 0.7750
qwen3-coder-plus Reasoning CPRetriever-Prob 0.7800
As shown, incorporating RAG improves Pass@1 across all evaluated models. Non-reasoning models benefit more noticeably, while reasoning models achieve smaller gains due to their strong baseline performance. Notably, using CPRetriever-Prob consistently outperforms Qwen3-Embedding-4B across all LLMs.
These results demonstrate that effective problem retrieval can enhance competitive programming problem-solving in a RAG setting, highlighting a valuable downstream application for our retrieval model.
A.4 Challenges in Competitive Programming Problem Retrieval
Retrieving competitive programming (CP) problems poses distinct challenges compared to standard code retrieval tasks in software engineering (SE) benchmarks such as SWE-Bench. The core difficulty lies in bridging the large semantic gap between a problem’s narrative description and its abstract algorithmic solution, whereas SE tasks typically involve more concrete, context-specific code retrieval.
CP problem statements are often highly abstract and indirect. They are framed as puzzles embedded within narratives (e.g., "a farmer needing a route") and require deep semantic understanding to map natural language to algorithmic concepts like graph traversal or dynamic programming. Critical information such as data constraints ($N\leq 1{,}000{,}000$) or time limits is subtly embedded, yet it fundamentally dictates the algorithmic complexity required ($O(N^{2})$ vs. $O(N\log N)$). Similarly, user queries are abstract (e.g., "longest increasing subsequence sum"), in contrast to SE queries, which are typically specific and grounded in concrete code modifications or bug fixes.
The mapping from a CP problem to its solutions is also more complex. A single problem may admit multiple fundamentally different correct algorithmic approaches (e.g., BFS/DFS vs. dynamic programming for tree diameter). CP retrieval aims to uncover the underlying idea or technique that can inspire a solution strategy, rather than locating an exact code snippet. In SE retrieval, the goal is generally to find specific functions, modules, or API calls that can be directly reused in a known context.
In summary, CP retrieval resembles a mathematician exploring theorems for similar logical structures while ignoring superficial details, focusing on abstract strategies. SE retrieval, in contrast, is more like a mechanic finding the right replacement part for a known machine. This higher-level abstraction and the need to translate between human language and algorithmic logic make competitive programming problem retrieval uniquely challenging.
A.5 Further Details on the CPRet-PCPCD
A.5.1 Language and Problem Format Statistics
@@
A.5.2 Programming Language Distribution
@@
A.6 Data Source Details
@@
A.6.1 Competitive Programming Problems and Codes Data(CPRet-PCPCD)
@@
A.6.2 Duplicate Programming Problem Pairs
@@
A.6.3 Simplified and Full Problem Description Pairs
@@
A.7 Extended Results: Model Performance across Similarity Bins
@@
A.8 Cases of Retrieval Tasks
A.8.1 Text-to-Code and Code-to-Code
@@
https://atcoder.jp/contests/arc162/tasks/arc162_a Problem Statement There are NN people, numbered from 11 to NN, who participated in a round-trip race between two points. The following information is recorded about this race. The outward times of any two people were different, and person $i$ $(1\leq i\leq N)$ had the $i$-th fastest outward time. The round-trip times (the sum of the outward and return times) of any two people were different, and person $i$ $(1\leq i\leq N)$ had the $P_{i}$-th fastest round-trip time. The person (or persons) with the fastest return time was awarded the fastest return award. Here, $P_{1},P_{2},\dots,P_{N}$ is a permutation of $1,2,\dots,N$. How many people could have received the fastest return award? There are $T$ test cases. Answer each of them. Constraints $1\leq T\leq 500$ $2\leq N\leq 10^{3}$ $P_{1},P_{2},\dots,P_{N}$ is a permutation of $1,2,\dots,N$ All input values are integers The sum of NN over all test cases does not exceed $10^{3}$ Input Format The input is given from standard input in the following format:
@@ A.8.2 Problem-to-Duplicate
https://codeforces.com/contest/1702/problem/G2 Problem: G2. Passable Paths (hard version) You are given a tree with $n$ vertices numbered from 11 to $n$. A tree is an undirected connected graph with n1n-1 edges and no cycles. You are also given $q$ queries. In each query, you are given a set of $k$ vertices. For each query, determine whether there exists a simple path in the tree that passes through all the given $k$ vertices. Input The first line contains an integer $n$ ($2\leq n\leq 2\times 10^{5}$) — the number of vertices in the tree. Each of the next n1n-1 lines contains two integers $u$ and $v$ ($1\leq u,v\leq n$) — the endpoints of an edge in the tree. The next line contains an integer $q$ ($1\leq q\leq 2\times 10^{5}$) — the number of queries. Each of the next $q$ lines describes a query: The first integer $k$ ($2\leq k\leq n$) — the number of vertices in the query. Followed by $k$ integers $v_{1},v_{2},\ldots,v_{k}$ ($1\leq v_{i}\leq n$) — the vertices in the query. Output For each query, print YES if there exists a simple path that passes through all the given $k$ vertices. Otherwise, print NO.
https://www.codechef.com/JULY21A/problems/KPATHQRY Problem: KPATHQRY — Path Queries on Trees You’re given a tree with NN vertices numbered from 11 to $N$. Your goal is to handle $Q$ queries. For each query, you are given $K$ nodes $v_{1},v_{2},\ldots,v_{K}$. Find whether there exists a simple path in the tree that covers all the given vertices. Input The first line contains a single integer $T$ — the number of test cases. For each test case: The first line contains a single integer NN — the number of vertices. Each of the following N1N-1 lines contains two integers $u$ and $v$ — an edge in the tree. The next line contains a single integer $Q$ — the number of queries. Each of the following $Q$ lines describes a query: * Starts with $K_{i}$ — the number of vertices in the query. * Followed by $K_{i}$ integers: $v_{1},v_{2},\ldots,v_{K_{i}}$. Output For each query, print "YES" if a simple path covering all the given nodes exists, otherwise print "NO". You may print each character of the string in uppercase or lowercase (for example, "yEs", "yes", "Yes", and "YES" will all be treated as identical). Constraints $1\leq T\leq 10$ $1\leq N\leq 10^{5}$ $1\leq u,v,v_{j}\leq N$ $1\leq Q\leq 10^{5}$ $1\leq K_{i}\leq N$ The edges form a valid tree. All vertices in a single query are distinct. The sum of NN over all test cases does not exceed $2\cdot 10^{5}$. The sum of $K_{i}$ over all queries in a single test case does not exceed $10^{5}$. Subtasks Subtask #1 (100 points): Original constraints.
A.8.3 Simplified-to-Full